a)      Your manager realizes that incentive programs are important to promote the efficiency and productivity of the workforce. The manager came up with the following wage packages to induce employees’ motivation. As a head of human resources, which scheme would you suggest and which program will motivate the employees more?

·        Wage scheme 1: Each employee will start at RM 17 per hour and will work eight hours per day.

·        Wage scheme 2: each employee will receive RM 8 per hour and one-tenth of 1 per cent of profits (expected profits are RM 90,000 per day if everyone puts out maximum effort).



1
Expert's answer
2021-11-01T12:48:33-0400

Wage scheme 1:

Each employee get RM 17 per hour

And work 8 hours per day

In single Day employee get = 17 * 8 = RM 136 extra as a incentive

Wage scheme 2:

Each employee get RM 8 per hour

And work 8 hours per day and 1/10 of 1 percent of profits

Expected profits are RM 90,000 per day

1% of RM 90000 = RM 900

1/10 th of 900 = 90

In single Day employee get = 8 * 8 = RM 64 extra as a incentive

In single day employee get = 64 + 90 = RM 154 extra as a incentive

Summary

Wage scheme 2 will motivate the employees more as compered to Wage scheme 1.